# Dissecting CLIP: Decomposition with a Schur Complement-based Approach

### v2

@@ Abstract
The use of CLIP embeddings to assess the alignment of samples produced by text-to-image generative models has been extensively explored in the literature. While the widely adopted CLIPScore, derived from the cosine similarity of text and image embeddings, effectively measures the relevance of a generated image, it does not quantify the diversity of images generated by a text-to-image model. In this work, we extend the application of CLIP embeddings to quantify and interpret the intrinsic diversity of text-to-image models, which is responsible for generating diverse images from similar text prompts. To achieve this, we propose a decomposition of the CLIP-based kernel covariance matrix of image data into text-based and non-text-based components. Using the Schur complement of the joint image-text kernel covariance matrix, we perform this decomposition and define the matrix-based entropy of the decomposed component as the Schur Complement Entropy (SCE) score, a measure of the intrinsic diversity of a text-to-image model based on data collected with varying text prompts. Additionally, we demonstrate the use of the Schur complement-based decomposition to nullify the influence of a given prompt in the CLIP embedding of an image, enabling focus or defocus of embeddings on specific objects or properties for downstream tasks. We present several numerical results that apply our Schur complement-based approach to evaluate text-to-image models and modify CLIP image embeddings. The codebase is available at https://github.com/aziksh-ospanov/CLIP-DISSECTION
Figure 1: An example of the CLIP embedding decomposition. The top row features images composed of two distinct concepts, such as "pumpkin" and "ostrich." In the bottom row, the top four ImageNet images with the highest CLIP scores are displayed after the specified concept has been removed from the embedding.
Figure 2: Comparison of model-driven (left) and prompt-driven (right) diversity in 200 images generated for similar "cat and dog" prompts (left) vs. 200 images generated for diverse "cat and dog" prompts (right). While unconditional metrics favor the right-side case, the SCE score isolates intrinsic model diversity, suggesting greater model diversity on the left.
@@ 1 Introduction
Text-based generative models, which generate samples guided by input text, have gained significant attention in computer vision research. In particular, text-to-image models, which create images based on input text, have found numerous applications and are widely used across various content creation tasks. Given the important role of text-to-image models in many AI applications, their training and evaluation have been extensively studied in recent years. A comprehensive evaluation of these models, addressing relevance, diversity, and fidelity, is essential to ensure their effectiveness and adaptability across different use cases.
@@
As the CLIP model [1] provides a joint embedding for text and image data, CLIP embeddings have been widely used to evaluate and interpret the performance of text-to-image models. By calculating the cosine similarity between the CLIP embeddings of text and image data, the CLIPScore [2] serves as a metric for measuring the alignment between the text and the generated image. While the CLIPScore and similar uses of CLIP embeddings focus on evaluating the fidelity and relevance of generated samples, these embeddings have not yet been applied to assess the diversity of data produced by text-to-image models. Existing evaluation frameworks typically address diversity scores for unconditional generative models, such as Recall [3, 4], Coverage [5], and Vendi [6], which assess image samples independently of the input text.
@@
However, the diversity of images generated by a text-to-image model depends both on the variety of input text data and on the model’s intrinsic diversity, which produces different images in response to similar prompts. Consequently, existing diversity metrics for unconditional generative models cannot differentiate between diversity arising from varied prompts and diversity contributed by the model itself. In this work, we leverage CLIP embeddings to propose a framework for quantifying and interpreting the intrinsic diversity of both text-to-image and image captioning (image-to-text) models. The primary objective of our approach is to decompose the CLIP image embedding into a text-based component, influenced by the input text, and a non-text-based component, arising from the model’s inherent randomness.
To achieve this goal, we employ a kernel similarity function and focus on the kernel covariance matrix for the generated (text, image) pairs. Using a Gaussian elimination-based approach in a proper kernel space, we decompose the image sub-covariance matrix, $C_{II}$, into the sum of the Schur complement component, $C_{II}-C_{IT}C_{TT}^{-1}C_{TI}$, which represents the model-driven diversity component, and the remainder term $C_{IT}C_{TT}^{-1}C_{TI}$, which captures the prompt-driven variety in the generated images. Leveraging the estimated kernel covariance matrices from generated data, we define the matrix-based entropy of the Schur complement component, as Schur Complement Entropy (SCE), as a measure of model-driven diversity in the generated images.
@@
To illustrate the effectiveness of the SCE score in distinguishing model-driven diversity from prompt-driven effects, we present a comparative example. Figure 2 compares the SCE-Image model diversity scores for samples generated by DALL-E 3 [7]. The left-side subfigure represents model-driven diversity, showing images of diverse-breed cats and dogs generated using similar prompts like “cat and dog read a text.” The right-side subfigure illustrates prompt-driven diversity, featuring same-breed cats and dogs across diverse-occasion prompts. While unconditional diversity metrics such as Vendi and RKE report higher diversity for the right-side case, the SCE score, designed to isolate intrinsic model diversity from prompt-driven effects, indicates greater diversity in the left-side case. This example highlights the advantage of the prompt-conditioned SCE score over existing unconditional diversity metrics.
@@
We present several numerical applications of the proposed framework to evaluate and interpret standard text-to-image and image-captioning models. Our results on several simulated scenarios with known ground-truth diversity indicate that the proposed entropy metric correlates with the non-text-based diversity in images, capturing variation not attributable to the text prompt. Additionally, we show that the decomposed feature component can neutralize the influence of specific objects in the text prompt within the image embedding. Specifically, we use this decomposed feature to diminish the impact of visible text in images, reducing its effect on the embedding. We also demonstrate how the decomposed image embedding can enhance or reduce the focus on particular objects or styles within an image, which can have implications for downstream applications of CLIP embeddings when emphasizing specific elements. For instance, Figure 1 shows a generated image of an ostrich next to a pumpkin. We demonstrate the ImageNet samples with highest CLIPScore to the original image embedding and to the modified image embedding when subtracting the Schur complement-method embedding after canceling texts "pumpkin" and "ostrich", which shows the modified embedding removes the effect of the unwanted object. The main contributions of this work are summarized as follows:
@@
  • Presenting numerical results on the Schur complement-based decomposition of CLIP embeddings
@@ 2 Related Work
CLIP interpretability and decomposition. Contrastive vision-and-language models, such as CLIP [1] are a class of models that were trained on paired text prompts and images. The notable feature of CLIP is a shared embedding space between image and text data. A common interpretability method involves heatmaps to highlight relevant image areas [8, 9, 10]. However, heatmaps are used to identify objects and lack spatially dependent information, such as object size and embedding output. Other approaches require decomposing the model architecture and analyzing attention heads. [11] introduced TextSpan, which finds a vector direction for each attention head and assigns it an appropriate text label. Another prominent approach, proposed in [12], decomposes dense CLIP embeddings into sparse, interpretable semantic concepts to enhance embedding description. [13] suggests disentangling written words from underlying concepts via orthogonal projection of learned CLIP representations.
Evaluation of Diversity in Generative Models. Diversity evaluation in generative models has been extensively explored in the literature, with existing metrics broadly classified as either reference-based or reference-free. Reference-based metrics, such as FID [14], KID [15], and IS [16], assess diversity by quantifying the distance between true and generated data. Metrics like Density and Coverage [5], and Precision and Recall [4] evaluate quality and diversity by analyzing the data manifold. Reference-free metrics, including Vendi [6, 17], RKE [18], and their variations like Nyström-Vendi and FKEA-Vendi [19], measure the entropy of a similarity metric (e.g., kernel matrix), capturing the number of distinct modes in the data. Since these metrics do not require a reference distribution, they are particularly suitable for text-to-image model evaluation, where selecting an appropriate reference dataset is challenging.
@@ 3 Preliminaries
Kernel Functions. We call $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ a kernel function if for samples $x_{1},\ldots,x_{n}\in\mathcal{X}$, the resultant kernel matrix $K=\bigl{[}k(x_{i},x_{j})\bigr{]}_{1\leq i,j\leq n}$ is a PSD (positive semi-definite) matrix. Moreover, $K$ can be decomposed into a dot product of feature maps $\phi$, referred to as the kernel trick, as follows:
@@
In our numerical analysis, we also consider the Gaussian Kernel defined as:
@@
Note that both of the presented kernels are normalized kernels, since the self-kernel similarity $k(x,x)=1$ for every $x\in\mathcal{X}$. In order to apply entropy measures, we consider the normalized kernel matrix given by $\frac{1}{n}K$. Therefore, we observe $Tr(\frac{1}{n}K)=1$ for every normalized kernel, implying that the eigenvalues of $\frac{1}{n}K$ form a probability measure, as they are non-negative and sum up to $1$.
@@
$H(A):=\mathrm{Tr}\bigl{(}-A\log A\bigr{)}=\sum_{i=1}^{n}\lambda_{i}\log\frac{1}{\lambda_{i}},$ (4)
where $\mathrm{Tr}(\cdot)$ denotes the matrix trace. As demonstrated in quantum information theory, the Von Neumann entropy measures the spread of eigenvalues. [6, 18] show that it can effectively count the number of sample clusters in the collected data. As noted by [6, 18], $\frac{1}{n}K$ shares the same non-zero eigenvalues with the kernel covariance matrix $\widehat{C}_{\mathcal{X}}$ defined as:
$\widehat{C}_{\mathcal{X}}:=\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\phi(x_{i})^{\top}=\frac{1}{n}\Phi^{\top}\Phi$ (5)
where $\Phi\in\mathbb{R}^{n\times d}$ is an $n\times d$ matrix whose rows are the feature presentations of samples $\phi(x_{1}),\ldots,\phi(x_{n})$. Feature representation varies between kernel methods. In Cosine Similarity Kernel, $\phi(x)=\text{CLIP}(x)/\|\text{CLIP}(x)\|_{2}$, i.e. normalized CLIP embedding of sample $x$, whereas in shift-invariant kernels, $\phi$ is a proxy feature map of the Gaussian kernel following the random Fourier features [26].
@@
$\Lambda=\begin{bmatrix}B&C\\ C^{T}&D\end{bmatrix}$
where $B$ and $D$ are square symmetric submatrices. If $B$ is invertible, the Schur Complement of $B$ in $A$ is given by $S=D-C^{T}B^{-1}C$. In general, even if $B$ is not invertible, $B^{-1}$ can be replaced with the Moore-Penrose psuedoinverse $B^{\dagger}$ in the Schur complement definition. Note that the Schur complement $S\succeq\mathbf{0}$ will be PSD for a PSD matrix $\Lambda\succeq\mathbf{0}$.
@@ 4 Diversity Evaluation for Text-to-Image Generative Models via CLIP Embedding
$C_{\text{\rm joint (I,T)}}:=\begin{bmatrix}C_{II}&C_{IT}\\ C_{IT}^{\top}&C_{TT}\end{bmatrix}$
@@
Proposition 1.
Define text-to-image conversion matrix $\Gamma^{*}=C_{TI}C_{TT}^{-1}$. Then, $\Gamma^{*}$ is an optimal solution to the following ($\|\cdot\|_{F}$ denotes the Frobenius norm):
@@
Then, in equation 6, $\Lambda_{T}$ is the covariance matrix of $\Gamma^{*}\phi(x_{T})$, and $\Lambda_{I}$ is the kernel covariance matrix of $\phi(x_{I})-\Gamma^{*}\phi(x_{T})$.
Remark 1.
Proposition 1 shows that given the optimal text-to-image conversion matrix $\Gamma^{*}=C_{TI}C_{TT}^{-1}$, the effect of a text $T$ on the embedding of an image $I$ can be canceled by considering the remainder term $\phi(x_{I})-\Gamma^{*}\phi(x_{T})$ to decorrelate the image and text embedding.
Following the above discussion for the cosine-similarity kernel $\phi(x)=x/\|x\|$, we can modify the CLIP embedding given a paired dataset of prompts and images. To do this, we first compute the modification matrix $\Gamma^{*}=C_{TI}C_{TT}^{-1}$ and modify the CLIP embedding for an image sample $x_{I}=\mathrm{CLIP}(I)$ and correcting prompt $x_{T}=\mathrm{CLIP}(T)$ as:
@@
Furthermore, in the decomposition in equation 6, both the image component $\Lambda_{I}$ and the text component $\Lambda_{T}$ are PSD matrices, which have non-unit trace values. To apply the matrix-based entropy definition which requires the unit trace, we rewrite the identity as
@@
Following the above decomposition, we define the Schur-Complement-Entropy for the image part $\mathrm{SCE_{I}}$ as follows:
$\mathrm{SCE}_{I}(x_{1},\ldots,x_{n})\,:=$ $\,\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}\log\frac{\mathrm{Tr}(\Lambda_{I})}{\lambda^{(\Lambda_{I})}_{j}}$ (8)
where $\lambda^{(\Lambda_{I})}_{j}$ denotes the $j$th eigenvalue of matrix $\Lambda_{I}$ and $\mathrm{Tr}(\Lambda_{I})=\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}$ is the sum of the eigenvalues. It can be seen that the defined Schur complement entropy is the conditional entropy of the hidden image cluster variable $\mathrm{Mode}\in\{1,2,\ldots,d\}$, which follows the spectral decomposition of the image covariance matrix $C_{II}=\sum_{j=1}^{d}\lambda_{j}v_{j}v_{j}^{\top}$, where the mode $\mathrm{Mode}=i$ has probability $\lambda_{i}$. Here we define a text-based guess of the image mode $Y_{T}$, which as we discuss in the Appendix can correctly predict the cluster with probability $\mathrm{Tr}(\Lambda_{T})$ and outputs erasure $e$ with probability $\mathrm{Tr}(\Lambda_{I})$. Then, we show in the Appendix that following standard Shannon entropy definition:
$\mathrm{SCE}_{I}(x_{1},\ldots,x_{n})=H\bigl{(}\mathrm{Mode}|Y_{T}\bigr{)}$

### v3

# Scendi Score: Prompt-Aware Diversity Evaluation via Schur Complement of CLIP Embeddings

@@ Abstract
The use of CLIP embeddings to assess the fidelity of samples produced by text-to-image generative models has been extensively explored in the literature. While the widely adopted CLIPScore, derived from the cosine similarity of text and image embeddings, effectively measures the alignment of a generated image, it does not quantify the diversity of images generated by a text-to-image model. In this work, we extend the application of CLIP embeddings to quantify and interpret the intrinsic diversity of text-to-image models, which are responsible for generating diverse images from similar text prompts, which we refer to as prompt-aware diversity. To achieve this, we propose a decomposition of the CLIP-based kernel covariance matrix of image data into text-based and non-text-based components. Using the Schur complement of the joint image-text kernel covariance matrix, we perform this decomposition and define the matrix-based entropy of the decomposed component as the Schur Complement ENtopy DIversity (Scendi) score, as a measure of the prompt-aware diversity for prompt-guided generative models. Additionally, we discuss the application of the Schur complement-based decomposition to nullify the influence of a given prompt on the CLIP embedding of an image, enabling focus or defocus of the embedded vectors on specific objects. We present several numerical results that apply our proposed Scendi score to evaluate text-to-image and LLM (text-to-text) models. Our numerical results indicate the success of the Scendi score in capturing the intrinsic diversity of prompt-guided generative models. The codebase is available at https://github.com/aziksh-ospanov/scendi-score.
{strip}[Uncaptioned image]
Figure 1: Comparison of model-driven diversity (left) and prompt-driven diversity (right) in 200 images generated for similar ”cat and dog” prompts (left) vs. 200 images generated for diverse ”cat and dog” prompts (right). The diversity metrics for unconditional generation, Vendi and RKE, favor the right-side case, but the Scendi score captures intrinsic model diversity, suggesting the left side has higher diversity.
@@ 1 Introduction
Prompt-guided generative models, which generate data guided by an input text prompt, have gained significant attention in the computer vision community. In particular, text-to-image and text-to-video models, which create visual data based on input text, have found many applications and are widely used across various content creation tasks. Given the important role of prompt-guided generative AI models in numerous machine learning applications, their training and evaluation have been extensively studied in recent years. A comprehensive evaluation of these models, addressing fidelity and diversity, is essential to ensure their effectiveness and adaptability across different use cases.
@@
As the CLIP model [45] provides a joint representation for text and image data, the CLIP embeddings have been widely used to evaluate and interpret the performance of text-to-image models. By calculating the cosine similarity between the CLIP embeddings of text and image data, the CLIPScore [14] serves as a fidelity metric for measuring the alignment between the text and the generated image. While the CLIPScore and similar uses of CLIP embeddings focus on evaluating the quality of generated samples, these embeddings have not yet been applied to assess the diversity of data produced by text-to-image models. Existing evaluation frameworks typically address diversity scores for unconditional sample generation (considering no input prompt), such as Recall [50, 28], Coverage [39], RKE [20], and Vendi [9], which assess image samples independently of the input text.
@@
However, the diversity of images generated by a text-to-image model depends both on the variety of input text data and on the model’s intrinsic diversity, driving the model to produce varied images in response to similar prompts. Therefore, existing diversity metrics for unconditional sample generation cannot differentiate between diversity arising from varied prompts, i.e., prompt-driven diversity, and diversity contributed by the model itself, i.e., model-driven diversity. In this work, we leverage CLIP embeddings to propose a framework for quantifying and interpreting the intrinsic diversity of both text-to-image and image captioning (image-to-text) models. The primary objective of our approach is to decompose the CLIP image embedding into a text-based component, influenced by the input text, and a non-text-based component, arising from the model’s inherent randomness.
To achieve this goal, we extend the kernel matrix entropy measures, i.e., Vendi and RKE scores, to enable a prompt aware diversity measurement. To do this, we consider a kernel similarity function and focus on the kernel covariance matrix for the generated (text, image) pairs. Considering the matrix-based entropy of the Image-based kernel covariance component $C_{II}$, one obtains the existing prompt-unaware Vendi [9] and RKE [20] scores. In this work, we propose applying the Gaussian elimination approach, and then decompose the image sub-covariance matrix, $C_{II}$ as follows:
$C_{II}=\underbrace{C_{II}-C_{IT}C_{TT}^{{}^{\text{\scriptsize$-1$}}}C^{{}^{\text{\scriptsize$\top$}}}_{IT}}_{\text{\rm Model-driven Covariance Component}}+\underbrace{C_{IT}\,C_{TT}^{{}^{\text{\scriptsize$-1$}}}\,C_{IT}^{{}^{\text{\scriptsize$\top$}}}}_{\text{\rm Prompt-driven Covariance Component}}$
Note that the above is the sum of the Schur Complement component, $C_{II}-C_{IT}C_{TT}^{-1}C_{IT}^{\top}$, which represents the model-driven diversity component, and the remainder term $C_{IT}C_{TT}^{-1}C_{TI}$, which captures the prompt-driven variety in the generated images. Extending the Vendi and RKE score to prompt-aware diversity evaluation, we define the matrix-based entropy of the Schur complement component, as Schur Complement ENtropy DIversity (Scendi) score. We highlight that Scendi is a measure of model-driven diversity in the prompt-guided generated data.
Figure 2: Example of the CLIP embedding decomposition for an input image of ”Pumpkin next to Ostrich” generated by SDXL [44]. ImageNet [7] samples with the highest CLIP similarity to the input image after removing the “pumpkin” direction from its CLIP embedding via the Schur complement (left) and after removing the “ostrich” direction (right).
@@
To illustrate the effectiveness of the Scendi score in distinguishing model-driven diversity from prompt-driven effects, we present a comparative example. Figure 1 compares the Scendi diversity scores for two sets of samples generated by DALL-E 3 [40]. The left-side subfigure is designed to represent model-driven diversity, showing images of diverse-breed cats and dogs generated using similar prompts like “cat and dog read a text.” In contrast, the right-side subfigure highlights the effects of prompt-driven diversity, featuring same-breed cats and dogs across diverse-occasion prompts. The unconditional diversity metrics, Vendi and RKE, report higher diversity for the right-side case. In contrast, our proposed Scendi score, designed to isolate intrinsic model diversity from prompt-driven effects, indicates greater diversity in the left-side case. This example highlights the advantage of the prompt-aware diversity measurement by Scendi score over existing unconditional diversity metrics in capturing the model-driven diversity.
@@
We present several numerical applications of the proposed framework to evaluate and interpret standard text-to-image and image-captioning models. Our results on several simulated scenarios with known ground-truth diversity indicate that the proposed entropy metric correlates with the non-text-based diversity in images, capturing variation not attributable to the text prompt. Additionally, we show that the decomposed feature component can neutralize the influence of specific objects in the text prompt within the image embedding. Specifically, we use this decomposed feature to diminish the impact of visible text in images, reducing its effect on the embedding. We also demonstrate how the decomposed image embedding can enhance or reduce the focus on particular objects or styles within an image, which can have implications for downstream applications of CLIP embeddings when emphasizing specific elements. For instance, Figure 2 shows a generated image of an ostrich next to a pumpkin. We showcase the ImageNet samples that achieve the highest CLIPScore when compared to the modified image embeddings obtained by canceling the “pumpkin” and “ostrich” directions. This demonstrates that the modified embeddings effectively eliminate the influence of those unwanted objects. The main contributions of this work are summarized as follows:
@@
  • Presenting numerical results on the Schur complement-based decomposition of CLIP embeddings and Scendi score performance under various diversity scenarios and data modalities.
@@ 2 Related Work
CLIP interpretability and decomposition. Contrastive vision-and-language models, such as CLIP [45] are a class of models that were trained on paired text prompts and images. The notable feature of CLIP is a shared embedding space between image and text data. A common interpretability method involves heatmaps to highlight relevant image areas [52, 53, 5, 11]. However, heatmaps are used to identify objects and lack spatially dependent information, such as object size and embedding output. Other approaches require decomposing the model architecture and analyzing attention heads. [10] introduced TextSpan, which finds a vector direction for each attention head and assigns it an appropriate text label. Another prominent approach, proposed in [2], decomposes dense CLIP embeddings into sparse, interpretable semantic concepts to enhance embedding description. [37] suggests disentangling written words from underlying concepts via orthogonal projection of learned CLIP representations. [4] utilize semantic guidance to move between concepts and introduce edits during diffusion process. Recently, [32] studied the geometry of the CLIP embeddings and identified numerous properties such as modality mass of image and text components in the embedding space.
Evaluation of Diversity in Generative Models. Diversity evaluation in generative models has been extensively explored in the literature, with existing metrics broadly classified as either reference-based or reference-free. Reference-based metrics, such as FID [15], KID [3], and IS [51], assess diversity by quantifying the distance between true and generated data. Metrics like Density and Coverage [39], and Precision and Recall [28] evaluate quality and diversity by analyzing the data manifold. Reference-free metrics, including Vendi [9, 43], RKE [20], and their improved and convergent variants FKEA-Vendi [42] and Truncated Vendi [41], measure the entropy of a similarity metric (e.g., kernel matrix), capturing the number of distinct modes in the data. Since these metrics do not require a reference distribution, they are particularly suitable for text-to-image model evaluation, where selecting an appropriate reference dataset is challenging. Another concurrent work on prompt-aware diversity is the Conditional Vendi score that measures the entropy of the Hadamard kernel product subtracted by the text component. [21]. Also, the kernel-based evaluation scores have been further utilized in the context of online model selection [16, 49, 17, 18], novelty evaluation and detection [55, 56], distributed evaluation [54], and embedding comparison and alignment [23, 12].
@@ 3 Preliminaries
Kernel Functions. We call $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ a kernel function if for samples $x_{1},\ldots,x_{n}\in\mathcal{X}$, the resultant kernel matrix $K=\bigl{[}k(x_{i},x_{j})\bigr{]}_{1\leq i,j\leq n}$ is a PSD (positive semi-definite) matrix. Moreover, $K$ can be decomposed into a dot product of the kernel feature maps $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ as follows:
@@
Also, we consider the Gaussian Kernel defined as:
@@
Note that both of the presented kernels are normalized kernels, since $k(x,x)=1$ holds for every $x\in\mathcal{X}$. To apply entropy measures, we consider the normalized kernel matrix given by $\frac{1}{n}K$. Therefore, we observe $\mathrm{Tr}(\frac{1}{n}K)=1$ for every normalized kernel, implying that the eigenvalues of $\frac{1}{n}K$ form a probability model, since they are non-negative and sum up to 111. Note that $\mathrm{Tr}(\cdot)$ denotes the matrix trace.
@@
$H(A):=\sum_{i=1}^{n}\lambda_{i}\log\frac{1}{\lambda_{i}}.$ (4)
[9] discusses that the Von-Neumann entropy of the normalized kernel matrix $\frac{1}{n}K$ can effectively capture the entropy of the cluster variable in the collected data, and proposes the Von Neumann entropy diversity (Vendi) score for measuring the diversity of a sampleset. Observe that $\frac{1}{n}K$ shares the same non-zero eigenvalues with the kernel covariance matrix $C_{X}$ defined as:
$C_{X}:=\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\phi(x_{i})^{\top}=\frac{1}{n}\Phi^{\top}\Phi$ (5)
where $\Phi\in\mathbb{R}^{n\times d}$ is an $n\times d$ matrix whose rows are the feature presentations of samples $\phi(x_{1}),\ldots,\phi(x_{n})$. Therefore, given the eigenvalues $\lambda_{1},\ldots,\lambda_{d}$ of the kernel covariance matrix $C_{X}$, the Vendi and RKE [20] scores are
$\mathrm{Vendi}(x_{1},\ldots,x_{n})$ $=\exp\Bigl{(}\sum_{i=1}^{d}\lambda_{i}\log\frac{1}{\lambda_{i}}\Bigr{)},$
$\mathrm{RKE}(x_{1},\ldots,x_{n})$ $=\frac{1}{\sum_{i=1}^{d}\lambda_{i}^{2}}=\frac{1}{\|C_{X}\|^{2}_{F}}.$
Feature representation varies between kernel methods. In Cosine Similarity Kernel, $\phi(x)=\text{CLIP}(x)/\|\text{CLIP}(x)\|_{2}$, i.e. normalized CLIP embedding of sample $x$, whereas in shift-invariant kernels, $\phi$ is a proxy feature map of the Gaussian kernel following the random Fourier features [46].
@@
$\Lambda=\begin{bmatrix}B&C\\ C^{\top}&D\end{bmatrix}$
where $B$ and $D$ are square symmetric submatrices. If $B$ is invertible, the Schur Complement of $B$ in $A$ is given by $S=D-C^{\top}B^{-1}C$. In general, even if $B$ is not invertible, $B^{-1}$ can be replaced with the Moore-Penrose psuedoinverse $B^{\dagger}$ in the Schur complement definition. Note that the Schur complement $S\succeq\mathbf{0}$ will be PSD for a PSD matrix $\Lambda\succeq\mathbf{0}$.
Figure 3: Evaluated $\mathrm{Scendi}$ scores with fixed text prompt distribution about doctors in various settings. Interpolated surface plot in the bottom right visualises the change in diversity (z-axis) when doctors have a varying diversity of age and gender features. (x and y axis)
@@ 4 Diversity Evaluation for Text-to-Image Generative Models via CLIP Embedding
$C_{\text{\rm joint (I,T)}}:=\begin{bmatrix}C_{II}&C_{IT}\vskip 5.69054pt\\ C_{IT}^{{}^{\text{\scriptsize$\top$}}}&C_{TT}\vskip 2.84526pt\end{bmatrix}$
@@
Proposition 1
Define text-to-image conversion matrix $\Gamma^{*}=C_{IT}C_{TT}^{-1}$. Then, $\Gamma^{*}$ is an optimal solution to the following ($\|\cdot\|_{F}$ denotes the Frobenius norm):
@@
Then, in (6), $\Lambda_{T}$ is the covariance matrix of $\Gamma^{*}\phi(x_{T})$, and $\Lambda_{I}$ is the kernel covariance matrix of $\phi(x_{I})-\Gamma^{*}\phi(x_{T})$.
Remark 1
Proposition 1 shows that given the optimal text-to-image conversion matrix $\Gamma^{*}=C_{IT}C_{TT}^{-1}$, the effect of a text $T$ on the embedding of an image $I$ can be canceled by considering the remainder term $\phi(x_{I})-\Gamma^{*}\phi(x_{T})$ to decorrelate the image and text embedding.
If we consider the cosine-similarity kernel feature map $\phi(x)=x/\|x\|$, the above discussion reveals an approach to cancel the effect of the input text on the CLIP embedding of the output images. Given a paired dataset of prompts and images, we first compute the modification matrix $\Gamma^{*}=C_{IT}C_{TT}^{-1}$ and modify the CLIP embedding for the input image $x_{I}=\mathrm{CLIP}(I)$ and prompt $x_{T}=\mathrm{CLIP}(T)$ as:
@@
Figure 4: Evaluated $\mathrm{Scendi}$, Vendi, Recall and Coverage scores with Gaussian Kernel on different cat breeds dataset.
Figure 5: Evaluated $\mathrm{Scendi}$ scores with fixed text prompt distribution. $\mathrm{Scendi}$ changes to reflect the diversity contributed by the varying level of information in text prompts.
Furthermore, in the decomposition in (6), both the model-driven covariance component $\Lambda_{I}$ and the text prompt-driven covariance component $\Lambda_{T}$ are PSD matrices, which have non-unit trace values. To apply the matrix-based entropy definition which requires the unit trace, we rewrite the identity as
@@
Definition 1
We define the Schur-Complement-ENtropy Diversity ($\mathrm{Scendi}$) score as follows:
$\mathrm{Scendi}(x_{1},..,x_{n};t_{1},..,t_{n})\,:=$ $\,\exp\Bigl{(}\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}\log\frac{\mathrm{Tr}(\Lambda_{I})}{\lambda^{(\Lambda_{I})}_{j}}\Bigr{)}$
where $\lambda^{(\Lambda_{I})}_{j}$ denotes the $j$th eigenvalue of matrix $\Lambda_{I}$ and $\mathrm{Tr}(\Lambda_{I})=\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}$ is the sum of the eigenvalues.
It can be seen that the defined Schur complement entropy is the conditional entropy of the hidden image cluster variable $\mathrm{Mode}\in\{1,2,\ldots,d\}$, which follows the spectral decomposition of the image covariance matrix $C_{II}=\sum_{j=1}^{d}\lambda_{j}v_{j}v_{j}^{\top}$, where the mode $\mathrm{Mode}=i$ has probability $\lambda_{i}$. Here we define a text-based guess of the image mode $Y_{T}$, which as we discuss in the Appendix can correctly predict the cluster with probability $\mathrm{Tr}(\Lambda_{T})$ and outputs erasure $e$ with probability $\mathrm{Tr}(\Lambda_{I})$. Then, we show in the Appendix that following standard Shannon entropy definition:
$\mathrm{Scendi}(x_{1},..,x_{n};t_{1},..,t_{n})=\exp\Bigl{(}H\bigl{(}\mathrm{Mode}|Y_{T}\bigr{)}\Bigr{)}$